OpenStudy (anonymous):

So how do i solve this. Answer this with explanation please or steps. 3i +2i^3 So how do i solve this. Answer this with explanation please or steps. 3i +2i^3 @Mathematics

OpenStudy (asnaseer):

\[2i^3=2i*i^2=2i*(-1)=-2i\]therefore: \[3i+2i^3=3i-2i=i\]

OpenStudy (asnaseer):

@JohnJungg - is this understandable or do you want more explanation?

OpenStudy (anonymous):

Is there anyway to simplify the 2i^3 in a different way?

OpenStudy (asnaseer):

since\[i=\sqrt{-1}\]\[i*i=-1\]therefore:\[2i^3=2*i*i*i=2*i*(-1)=-2i\]

OpenStudy (anonymous):

Alright thank you

OpenStudy (asnaseer):

yw